# Adapt PointFormer: 3D Point Cloud Analysis via Adapting 2D Visual Transformers

### v1

@@ 1 Introduction
We first devise an empirical study to further investigate the viability of directly applying image priors in point cloud analysis. A randomly initialized lightweight PointNet (RPN) is utilized for aligning the dimensions of point clouds with those of image tokens and then obtain the random point embedding. The dimension alignment network is fixed during training. Subsequently, a pre-trained 2D transformer undergoes fine-tuning with the sequenced random point embedding as input. The results are shown in Figure 1. It can be observed that, compared to the model trained on 3D data from scratch, the fine-tuned 2D model attains higher accuracy. Therefore, the attention derived from pre-trained models on 2D images exhibits efficacy in analyzing 3D point clouds.
@@ 2 Related work
Self-Attention-based Methods. Self-attention operations [43] have been adopted for point cloud processing in several studies [55, 13, 7]. For example, the point transformer [55] and point cloud transformer (PCT) [13] have introduced self-attention networks [43] to improve the capture of local context within the point clouds. Afterward, a plethora of methods based on the self-attention architecture have been proposed. PointMixer [7] enhances self-attention layers through inter-set and hierarchical-set mixing. TokenFusion [46] initially fuses tokens from heterogeneous modalities with point clouds and images, subsequently forwarding the fused tokens to a shared transformer, allowing learning of correlations among multimodal features. AShapeFormer [25] utilizes multi-head attention to effectively encode information pertaining to object shapes. This encoding capability can be seamlessly integrated with established 3D object detection methodologies. Exploiting pre-trained transformer models is also a promising way. P2P [47] employs a lightweight DGCNN [45] for the conversion of point clouds into visually rich and informative images, which serves to facilitate the utilization of pre-trained 2D knowledge. Point-BERT [52] constructs point cloud tokens that represent various geometric patterns, resembling word tokens. Subsequently, pre-trained language models, represented by BERT [8], can be applied to downstream tasks such as object classification, part segmentation, and related applications.
@@ 2.2 Parameter-Efficient Fine-Tuning
Recent advancements for 2D visual recognition incorporate the pre-trained Transformer models, exemplified by models like CLIP [36] and ViT [11]. Parameter-efficient fine-tuning (PEFT) [51], similar to model reprogramming or adversarial reprogramming [12] in the field of adversarial learning is designed to capitalize on the representational capabilities inherited from pre-trained models. PEFT strategically fine-tunes only a few parameters to achieve better performance in diverse downstream tasks that are different from the pre-trained models [4]. Representative methods, including prompt tuning (PT) [22], adapters [18], and Low-rank adapter (LoRA) [19] were initially designed for the purpose of incorporating language instructions into the input text for language models. LoRA [19] applies parameter tuning within the multi-head self-attention module in transformers. He et al. [16] introduced adapters into the field of computer vision. Bahng et al. [3] first defined the “visual prompt", mirroring the “prompt" in NLP. Subsequently, PEFT in the realm of computer vision, such as visual prompt tuning (VPT) [20], visual adapters [5, 30], and visual LoRA [38], to name a few, exhibit outstanding performance with minimal training parameters, reduced epochs, and substantial performance enhancements.
@@ 3.1 Motivation
Self-attention-based architecture, commonly referred to as transformer, has demonstrated noteworthy advancements in the analysis of point clouds [13, 55]. Nonetheless, the transformer architecture [43, 8, 11] exhibits inferior performance in comparison to CNNs when trained from scratch on a midsize dataset like imageNet-1K [53]. In contrast to more readily accessible 2D image data, the acquisition and annotation of 3D point cloud data imposes considerable financial and temporal burdens [27] due to its irregular and inhomogeneous character. This disparity poses a challenge in training transformer-based networks for point cloud analysis. We, therefore, propose to utilize PEFT technology, which demands less training data. However, existing pre-training models predominantly rely on 2D image data. To effectively utilize these pre-trained models, the calibration of dimensions and attention for 3D point clouds with 2D image prior becomes crucial.
@@ 3.2 Preliminaries
2D Visual Transformer. A visual transformer (ViT) model comprises an embedding layer and multiple transformer blocks. For an input image $x^{I}$, the model first partitions $x^{I}$ into $m$ patches, forming a set $\{x^{I}_{i}\}_{i=1}^{m}$. These patches are then embedded into sequences of $d^{I}$-dimensional vectors, denoted as $E^{I}_{0}=\texttt{Embed}\left(\left[x^{I}_{1},x^{I}_{2},\cdots,x^{I}_{m}\right]\right)$, where $E^{I}_{0}\in\mathbb{R}^{m\times d}$. $E^{I}_{0}$ is subsequently fed into $L$ blocks $\{\phi^{(l)}\}_{i=1}^{L}$ within the transformer model. We use the superscript $(l)$ to denote the index of the block. Formally, this procedural description can be mathematically expressed as:
@@
A singular self-attention within $\texttt{MSA}^{l}$ is calculated by softmax-weighted interactions among the input query, key, and value tokens obtained by three different learnable linear projection weights. Finally, the class prediction is achieved by a linear classification head.
@@ 3.4 PointFormer
The transformer-based architecture is more data-hungry than CNN-based ones [53]. In comparison to image data, the availability of 3D data is relatively constrained, resulting in issues such as overfitting and a limited realization of the transformer-based model potential. This paper investigates parameter-efficient fine-tuning (PEFT) technology to alleviate overfitting and improve model generalization for 3D models. PEFT involves the freezing of the pre-trained backbone that is previously trained on an extensive dataset, while introducing a limited number of learnable parameters to adapt to the new dataset. This new dataset can be data-rich [20, 3], few-shot [21], or long-tailed [9], as PEFT equips the model with knowledgeable priors. AdapterFormer [5] is an effective PEFT method. It appends the MLP layer in Eq (3) with a bottleneck module and has been empirically validated for its efficacy in handling 2D image data. We utilize this architecture for calibrating the point tokens alongside the 2D image attention, namely introducing a trainable bottleneck module. Formally, the calibration of point embeddings is calculated as follows:
@@ 3.5 Downstream Tasks
The principal disparity between APF and existing methods lies in that APF demonstrates the viability of adapting 2D priors to 3D feature space with minimal training parameters, rather than relying on specific network architectures. APF essentially executes “2D alignment to 3D”. The Morton order in the sequencing step aims to mimic the ordered image tokens fed in ViT. The PointFormer module facilitates the adaptation of prior attention in pre-trained ViT. Most existing methods perform “3D alignment to 2D”. For example, I2P-MAE [54] leverages 2D knowledge by projecting 3D point clouds onto multiple corresponding 2D images. P2P [47] transforms the 3D point cloud of an object into a single RGB image.
@@ 4.3 Further Analysis
The Impact of 2D Image Prior. We design an experiment to evaluate the impact of 2D priors on the 3D point cloud analysis. We employ ViT-B, pre-trained on ImageNet-21k, to furnish the 2D prior. A PointNet for input embedding is randomly initialized and, subsequently, its parameters are frozen, leaving only the PointFormer in APF as learnable. Under this setting (APF w. RPN in Table 6), the object classification accuracies are 92.2% and 80.1% on ModelNet40 and ScanObjectNN, respectively. APF exhibits considerable performance gains over PointNet by only utilizing randomly initialized embedding projection. Furthermore, it even outperforms the Transformer trained from scratch while incurring significantly lower training costs compared to training from scratch. This observation shows that the prior knowledge embedded in the 2D pre-trained model can significantly aid 3D point cloud analysis, even in scenarios where their training sets and data modalities differ. Moreover, the PointFormer proves beneficial in calibrating 2D-3D attention. Involvement of the point cloud projection in training, as indicated by APF with TPN in Table 6, further improves the performance of APF. This demonstrates the essential role of a representative point embedding in fully leveraging the potential offered by 2D prior knowledge.
Technical details for Morton-order. Sequencing is achieved by Morton code [29]. In detail, the process begins by selecting one point embedding as the initial point. Next, the coordinates of the center points corresponding to point embeddings are encoded into one-dimensional space using Morton code, and subsequently sorted to determine their order. This Morton order, also called Z-order ensures that point embeddings from the closest coordinates are adjacent. Figure 5 shows an example of the comparison between ordered and disordered point clouds. We randomly select 20 points for clear visualization. Figure 6 shows the Z-order sorting in 3D space.

### v2

@@ 1 Introduction
We first devise an empirical study to further investigate the viability of directly applying image priors in point cloud analysis. A randomly initialized lightweight PointNet (RPN) is utilized for aligning the dimensions of point clouds with that of image tokens and then obtaining the random point embedding. The dimension alignment network is fixed during training. Subsequently, a pre-trained 2D transformer undergoes fine-tuning with the sequenced random point embedding as input. The results are shown in Figure 1. It can be observed that, compared to the model trained on 3D data from scratch, the fine-tuned 2D model attains higher accuracy. Therefore, the attention derived from pre-trained models on 2D images exhibits efficacy in analyzing 3D point clouds.
@@ 2 Related work
Self-Attention-based Methods. Self-attention operations [45] have been adopted for point cloud processing in several studies [58, 13, 7]. For example, the point transformer [58] and point cloud transformer (PCT) [13] have introduced self-attention networks [45] to improve the capture of local context within the point clouds. Afterward, a plethora of methods based on the self-attention architecture have been proposed. PointMixer [7] enhances self-attention layers through inter-set and hierarchical-set mixing. TokenFusion [48] initially fuses tokens from heterogeneous modalities with point clouds and images, subsequently forwarding the fused tokens to a shared transformer, allowing the learning of correlations among multimodal features. AShapeFormer [25] utilizes multi-head attention to effectively encode information pertaining to object shapes. This encoding capability can be seamlessly integrated with established 3D object detection methodologies. Exploiting pre-trained transformer models is also a promising way. P2P [49] employs a lightweight DGCNN [47] for the conversion of point clouds into visually rich and informative images, which serves to facilitate the utilization of pre-trained 2D knowledge. Point-BERT [54] constructs point cloud tokens that represent various geometric patterns, resembling word tokens. Subsequently, pre-trained language models, represented by BERT [8], can be applied to downstream tasks such as object classification, part segmentation, and related applications.
@@ 2.2 Parameter-Efficient Fine-Tuning
Recent advancements for 2D visual recognition incorporate the pre-trained Transformer models, exemplified by models like CLIP [37] and ViT [11]. Parameter-efficient fine-tuning (PEFT) [53], similar to model reprogramming or adversarial reprogramming [12] in the field of adversarial learning is designed to capitalize on the representational capabilities inherited from pre-trained models. PEFT strategically fine-tunes only a few parameters to achieve better performance in various downstream tasks that are different from the pre-trained models [4]. Representative methods, including prompt tuning (PT) [22], adapters [18], and Low-rank adapter (LoRA) [19] were initially designed for the purpose of incorporating language instructions into the input text for language models. LoRA [19] applies parameter tuning within the multi-head self-attention module in transformers. He et al. [16] introduced adapters into the field of computer vision. Bahng et al. [3] first defined the “visual prompt", mirroring the “prompt" in NLP. Subsequently, PEFT in the realm of computer vision, such as visual prompt tuning (VPT) [20], visual adapters [5, 31], and visual LoRA [39], to name a few, exhibit outstanding performance with minimal training parameters, reduced epochs, and substantial performance enhancements. Recently, there have been proposed numerous PEFT methods for fine-tuning 3D pre-trained models [56, 41, 59]. Any2Point [26] proposes to utilize a PEFT method for fine-tuning pre-trained models with any modality, leveraging prior spatial knowledge. This approach demonstrates the efficacy of pre-trained knowledge of different modalities in enhancing 3D understanding.
@@ 3.1 Motivation
Self-attention-based architecture, commonly referred to as transformer, has demonstrated noteworthy advancements in the analysis of point clouds [13, 58]. Nonetheless, the transformer architecture [45, 8, 11] exhibits inferior performance in comparison to CNNs when trained from scratch on a midsize dataset like imageNet-1K [55]. In contrast to the more readily accessible 2D image data, the acquisition and annotation of 3D point cloud data imposes considerable financial and temporal burdens [28] due to its irregular and inhomogeneous character. This disparity poses a challenge in training transformer-based networks for point cloud analysis. We, therefore, propose to utilize PEFT technology, which demands less training data. However, existing pre-training models predominantly rely on 2D image data. To effectively utilize these pre-trained models, the calibration of dimensions and attention for 3D point clouds with 2D image prior becomes crucial.
@@ 3.2 Preliminaries
2D Visual Transformer. A visual transformer (ViT) model comprises an embedding layer and multiple transformer blocks. For an input image $x^{I}$, the model first partitions $x^{I}$ into $m$ patches, forming a set $\{x^{I}_{i}\}_{i=1}^{m}$. These patches are then embedded into sequences of $d^{I}$-dimensional vectors, denoted as $E^{I}_{0}=\texttt{Embed}\left(\left[x^{I}_{1},x^{I}_{2},\cdots,x^{I}_{m}\right]\right)$, where $E^{I}_{0}\in\mathbb{R}^{m\times d}$. $E^{I}_{0}$ is subsequently fed into $L$ blocks $\{\phi^{(l)}\}_{i=1}^{L}$ within the transformer model. We use superscript $(l)$ to denote the index of the block. Formally, this procedural description can be mathematically expressed as:
@@
A singular self-attention within $\texttt{MSA}^{l}$ is calculated by the softmax-weighted interactions among the input query, key, and value tokens obtained by three different learnable linear projection weights. Finally, the class prediction is achieved by a linear classification head.
@@ 3.4 PointFormer
The transformer-based architecture is more data-hungry than CNN-based ones [55]. In comparison to image data, the availability of 3D data is relatively constrained, resulting in issues such as overfitting and a limited realization of the transformer-based model potential. This paper investigates parameter-efficient fine-tuning (PEFT) technology to alleviate overfitting and improve model generalization for 3D models. PEFT involves the freezing of the pre-trained backbone that is previously trained on an extensive dataset, while introducing a limited number of learnable parameters to adapt to the new dataset. This new dataset can be data-rich [20, 3], few-shot [21], or long-tailed [9], as PEFT equips the model with knowledgeable priors. AdapterFormer [5] is an effective PEFT method. It appends the MLP layer in Eq (3) with a bottleneck module and has been empirically validated for its efficacy in handling 2D image data. We utilize this architecture for calibrating the point tokens alongside the 2D image attention, namely introducing a trainable bottleneck module. Formally, the calibration of point embedding is calculated as follows:
@@ 3.5 Downstream Tasks
The principal disparity between APF and existing methods lies in that APF demonstrates the viability of adapting 2D priors to 3D feature space with minimal training parameters, rather than relying on specific network architectures. APF essentially executes “2D alignment to 3D”. The Morton order in the sequencing step aims to mimic the ordered image tokens fed in ViT. The PointFormer module facilitates the adaption of prior attention in pre-trained ViT. Most existing methods perform “3D alignment to 2D”. For example, I2P-MAE [57] leverages 2D knowledge by projecting 3D point clouds into multiple corresponding 2D images. P2P [49] transforms the 3D point cloud of an object into a single RGB image.
@@ 4.3 Further Analysis
The Impact of 2D Image Prior. We design an experiment to evaluate the impact of 2D priors on the 3D point cloud analysis. We employ ViT-B, pre-trained on ImageNet-21k, to furnish the 2D prior. A PointNet for input embedding is randomly initialized and subsequently, its parameters are frozen, leaving only the PointFormer in APF as learnable. Under this setting (APF w. RPN in Table 6), the object classification accuracies are 92.2% and 80.1% on ModelNet40 and ScanObjectNN, respectively. APF exhibits considerable performance gains over PointNet by only utilizing randomly initialized embedding projection. Furthermore, it even outperforms the Transformer trained from scratch while incurring significantly lower training costs compared to training from scratch. This observation shows that the prior knowledge embedded in the 2D pre-trained model can significantly aid 3D point cloud analysis, even in scenarios where their training sets and data modalities differ. Moreover, the PointFormer proves beneficial in calibrating 2D-3D attention. Involvement of the point cloud projection in training, as indicated by APF with TPN in Table 6, further improves the performance of APF. This demonstrates the essential role of a representative point embedding in fully leveraging the potential offered by 2D prior knowledge.
Technical details for Morton-order. Sequencing is achieved by Morton code [30]. In detail, the process begins by selecting one point embedding as the initial point. Next, the coordinates of the center points corresponding to point embeddings are encoded into one-dimensional space using Morton code, and subsequently sorted to determine their order. This Morton order, also called Z-order ensures that point embeddings from the closest coordinates are adjacent. Figure 5 shows an example of the comparison between ordered and disordered point clouds. We randomly select 20 points for clear visualization. Figure 6 shows the Z-order sorting based on Z-value in 3D space.